Wilmington Company has two manufacturing departments—Assembly and Fabrication. It considers all of its manufacturing overhead costs to be fixed costs. The first set of data that is shown below is based on estimates from the beginning of the year. The second set of data relates to one particular job completed during the year—Job Bravo. Estimated Data Assembly Fabrication Total Manufacturing overhead costs $ 1,920,000 $ 2,240,000 $ 4,160,000 Direct labor-hours 80,000 48,000 128,000 Machine-hours 32,000 160,000 192,000 Job Bravo Assembly Fabrication Total Direct labor-hours 17 9 26 Machine-hours 9 12 21 Required: 1. If Wilmington used a plantwide predetermined overhead rate based on direct labor-hours, how much manufacturing overhead would be applied to Job Bravo? 2. If Wilmington uses departme ntal predetermined overhead rates with direct labor-hours as the allocation base in Assembly and machine-hours as the allocation base in Fabrication, how much manufacturing overhead would be applied to Job Bravo? (Round your intermediate calculation to 2 decimal places.)
Answer #1

1) Plantwide overhead rate = 4160000/128000 = 32.50 per DLH

Overhead applied to Job Bravo = 26*32.50 = 845

2) Assembly overhead rate = 1920000/80000 = 24 per DLH

Fabrication overhead rate = 2240000/160000 = 14 per MH

Overhead applied to Job Bravo = (17*24+12*14) = 576